the radius of cation is smaller than the parental atom .give reason
 

Grade:11

3 Answers

Mukhtar Hussain
6 years ago
Due to Loss of Electron, Shrinkage of electron Cloud, Decrease in Inter_Electronic Repulsion, Sometimes Loss of Shell
Chhavi Jain
6 years ago
The size of cation is smaller than the parent atom because cation is formed when the parent atom loose elections. Due to which the nuclear force of attraction ( Z effective ) increases on electrons so atom shrinks. Eg. SodiumNa ( p=11; e=11; p/e =11/ 11=1 )Na + ( p=11; e=10; p/e = 11/10=1.1) therefore in case of Na+ ion the nuclear force of attraction per electron increases due to which size decreases.
Sudhir Bishnoi
6 years ago
A cation forms when parent atom looses electrons while no. of protons in nucleus remains same. Therefore effective nuclear charge increases i.e. attraction between nucleus and electron increases and hence distance between them decreases i.e. radius decreases.